Dublin University Rifle Club

Version Date : April 23, 2002

DURC Competition Attendance Report

Fassaroe Open 2001

Distribution

Dr. Trevor West DUCAC Chairman
Cathy Doyle DUCAC Administrative Officer
Geraldine McAuley DUCAC Administrative Officer
Terry McAuley DURC President and Director of Sport
Michelle Bennett Physical Recreation Officer
Helen Murphy DUCAC Sports Facilities Committee Secretary

Summary

Venue Fassaroe Shooting Club, Enniskerry
Date June 16-17, 2001
Discipline 50m Three Position Rifle,

50m Prone Rifle, 100yd Prone Rifle DURC Attendees

Steven Watterson Mark Dennehy
John Keeney  Alan Smyth
Joe Lalor  Peter Byrne

Ray Reilly

Scores 50m Prone Rifle

Ray Reilly 571
Steven Watterson 570
Mark Dennehy 556
Peter Byrne 549
Joe Lalor 534

100yd Prone Rifle

John Keeney 555
Alan Smyth 511
Joe Lalor 338

Prizes Won

John Keeney First Place, 100yd Prone Rifle

Report

DURC attended the Fassaroe Open this weekend despite inclement weather and various other logistical difficulties. While we had no entrants for the Three Position match (which is the most rigorous discipline in target shooting and which cannot be trained for in Trinity due to range restrictions; it is an Olympic discipline), we had great success in the Scottish Match (100yd Prone Rifle), with John Keeney winning not only the handicapped section but the open section as well, with an excellent score of 555 – beating Olympic veterans Alan Lewis and Rhona Barry. The most heavily attended competition was, as usual, the English Match (50m Prone Rifle, which is an Olympic discipline also – the name is purely historical) and DURC acquitted themselves reasonably well here despite low scores from Mark Dennehy (who is currently breaking in a full set of new equipment). Joe Lalor, on his first outdoor shoot, did quite well, as did Peter Byrne. Steven Watterson shot somewhat below his normal standard, as did Ray Reilly, but hopefully some more training will rectify this now that exam season is completed. Overall, a satisfactory outdoor shoot given the conditions involved, but we hope to have more attendants and higher scores for the National 50m Championships at the beginning of July.

Misé le meas,

Mark Dennehy, DURC Secretary